OpenStudy (anonymous):

what are the zeros for x^2+2x+10

OpenStudy (anonymous):

just factor

OpenStudy (anonymous):

@LagrangeSon678 It does not factor out evenly

OpenStudy (anonymous):

Are you not using the quadratic formula yet?

OpenStudy (radar):

\[x ^{2}+2x+10=0\]\[x ^{2}+2x+1=-9\]\[(x+1)^{2}=-9\]\[x+1=\sqrt{-9}\]\[x=\sqrt{-9}-1\]\[x=\sqrt{9}\sqrt{-1}-1\]\[x=\pm3i-1\]

OpenStudy (radar):

Two complex roots

OpenStudy (anonymous):

@radar Thank You soo much

OpenStudy (radar):

u r most welcome and good luck